OpenStudy (x__hazel__x):

29. Nancy walked 516 yards in 6 minutes and then another 368 yards in 4 minutes. What was Nancy's average speed? A. 88.4 yd/min B. 89 yd/min C. 178 yd/min D. 89.4 yd/min

OpenStudy (x__hazel__x):

Is it similar to the last one?

TheSmartOne (thesmartone):

add total distance and divide it by the total time average speed = (total distance)/(total time)

OpenStudy (anwaarullah):

its same like last question

OpenStudy (anwaarullah):

so try

OpenStudy (x__hazel__x):

Okay, 516/6 = 86

OpenStudy (x__hazel__x):

Is that correct?

jimthompson5910 (jim_thompson5910):

hint: she walked 516+368 = 884 yards in 6+4 = 10 minutes

OpenStudy (anwaarullah):

nope take averg of these two thren

OpenStudy (x__hazel__x):

But Im trying to find the average

TheSmartOne (thesmartone):

so you have to find the total distance and the total time before using the formula :)

OpenStudy (x__hazel__x):

ohh, how do I do that?

OpenStudy (anwaarullah):

516/6= $$ and 368/4=&& so $$+&&/2

OpenStudy (x__hazel__x):

516/6 = 68 and 368/4 = 92

TheSmartOne (thesmartone):

Hazel, Jim gave you the total distance and total time up above :P

OpenStudy (x__hazel__x):

OHH I didnt see that, sorry XD

OpenStudy (x__hazel__x):

Thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (x__hazel__x):

Okay, so 884 yards in 10 minutes

OpenStudy (anwaarullah):

B option tho lol

OpenStudy (x__hazel__x):

lool

OpenStudy (x__hazel__x):

So Ill put A

OpenStudy (anwaarullah):

89 is correct :PP

OpenStudy (anwaarullah):

but you have to try yourself otherwise you wont get math, if you do wana excel in it :PP

OpenStudy (x__hazel__x):

Okay! Thank you!

OpenStudy (anwaarullah):

weclm